# On uniqueness of KP soliton structures

### v1

@@ Abstract.
G. C. was partially funded by Simons fundation MP-TSM-00002258.
@@ 1.1. Setting of the problem
where $u=u\left(t,x,y\right)\in\mathbb{R}$ is the unknown. The KP equations are canonical integrable models in 2D and were first introduced by Kadomtsev and Petviashvili in 1970 [14] for modeling “long and weakly nonlinear waves” propagating essentially along the $x$ direction, but with a small dependence on the $y$ variable. A rigorous derivation of both models from the Boussinesq system was obtained by Lannes and Lannes-Saut [24, 25].
@@
    The third line has a complex structure represented by a Heat type term defined by $H\left(\Theta\right):=\Theta_{y}-\Theta_{xx}$. Notice that $H$ is an operator in the $y$ and $x$ variables only.
@@
Later we will prove that under $H(\Theta)=0$, one has $W_{x}^{F}(\Theta)-W_{y}^{F}(\Theta)=W_{x}(\Theta)-W_{y}(\Theta)=0$, namely one can assume that $F=\log$ in (1.11), leading to the natural definitions in (1.6). In that sense, null Wronskians are naturally related to the Heat condition $H\left(\Theta\right)=0$, however, the equivalence will not be as exact as one would prefer.
@@ 1.2. KP solitons
Now we recall the KP 2-soliton. In this case $\Theta=\hbox{Wr}\left[\Theta_{1},\Theta_{2}\right]$, where Wr is the Wronskian of two functions $\Theta_{1}=\exp\left(\theta_{1}\right)+\exp\left(\theta_{2}\right)$ and $\Theta_{2}=\exp\left(\theta_{3}\right)+\exp\left(\theta_{4}\right)$ being 1-soliton phases. Calculating the phase $\Theta$, one obtains the classical formula
@@ 1.3. Main Results
This paper represents a departure from previous approaches, in a sense to be explained below. We adopt the perspective that each KP soliton should adhere to a specific “variational” characterization, manifested in their critical points of a suitable nonlinear functional. Our primary aim is to offer clear and straightforward characterizations of the most distinct KP solitons using simple ’trapping’ functionals that exhibit rigidity properties. This represents an initial step in the direction previously outlined. We believe that this concept holds promising potential for applications not only in elucidating more complex KP solutions but also in other related dispersive models.
@@
Let $u$ be a smooth solution to KP (1.1) of the form (1.2), with a smooth profile $F$ such that $F\left(1\right)=0$, $F^{\prime}\left(1\right)=1$, $F^{\prime\prime}\left(1\right)=-1$, and $F^{\prime\prime\prime}\left(1\right)=2$. Then $u$ is a KdV soliton and $F=\log$ if and only if $H\left(\Theta\right)=Ai\left(\Theta\right)=W_{y}^{F}\left(\Theta\right)=0.$
@@
Assume (1.17). Let $u$ be a nontrivial KP solution of the form (1.2) such that $\Theta>0$ is a solution to $H\left(\Theta\right)=Ai\left(\Theta\right)=W_{y}^{F}\left(\Theta\right)=0.$ Then $F=\log$ and $u=Q_{k}$ in (1.14) for some $k\in\mathbb{R}-\{0\}$.
@@
It is worth to mention that the conditions $H\left(\Theta\right)=Ai\left(\Theta\right)=0$ do not ensure the validity of Theorem 1.8. Indeed, it will be proved that the class of phases satisfying these two conditions is large enough to contain many multi-soliton solutions, such as the $Y$ structure defined below, which is not a line soliton. Consequently, (1.18) is a necessary condition.
@@
Theorem 1.8 puts in evidence an intriguing new mathematical structure, a natural finite-dimensional cone in the $y$ variable.
@@
The space $\mathcal{W}_{n}$ has interesting properties, in particular its behavior under the nonlinear mapping $\Theta W_{y}\left(\Theta\right)$ is key to understand complex multi-soliton structures. First of all, $\Theta\in\mathcal{W}_{M}$ implies that $\Theta W_{y}\left(\Theta\right)\in\mathcal{W}_{\frac{1}{2}M\left(M-1\right)}$ (Lemma 4.7). Second, there is a natural “kernel” given by the function $\exp\left(ky\right)$, $k$ arbitrary: one has $\exp\left(ky\right)W_{y}\left(\exp\left(ky\right)\right)=0$ for any $k=k\left(t,x\right)$. Additionally, under the gauge $\mathcal{W}_{M}\ni\Theta\longmapsto\Theta_{k}:=\exp\left(ky\right)\Theta$, one has $\Theta_{k}W_{y}\left(\Theta_{k}\right)=\exp\left(2ky\right)\Theta W_{y}\left(\Theta\right)\in\mathcal{W}_{\frac{1}{2}M(M-1)}$, revealing that there is no unique nontrivial solution to the set inclusion $\Theta W_{y}\left(\Theta\right)\in\mathcal{W}_{\frac{1}{2}M\left(M-1\right)}$ for $\Theta\in\mathcal{W}_{M}$. One way to repair this gauge freedom is to ask for $\Theta\left(y=0\right)$ and $\Theta_{y}\left(y=0\right)$ uniquely defined, as it is done below. [^1] On the other hand, it is easy to check that $\mathcal{W}:=\bigcup_{n\geq 0}\mathcal{W}_{n}$ is a multiplicative algebra under nonnegative coefficients. With respect to the $y$ variable, standard topological arguments ensure that the closure of this space restricted to any compact set $K$ of $\mathbb{R}$, under the uniform norm, is nothing but $C_{+}\left(K,\mathbb{R}\right)$, the space of real-valued, nonnegative continuous functions defined on $K$. For natural reasons, we will work in a slightly larger class of resonant phases than (1.15), given by
$\Theta\left(t,x,y\right)=\sum^{M}_{i=1}\left(a_{i,1}\exp\left(-k_{i}x+k^{2}_{i}y+k^{3}_{i}t\right)+a_{i,2}\exp\left(k_{i}x+k^{2}_{i}y+k^{3}_{i}t\right)\right),$ (1.20)
with coefficients $k_{1}<k_{2}<\cdots<k_{M}$ and $a_{i,j}>0$.
@@
Assume that the smooth profile $F$ such that (1.17) holds. Then $\Theta$ corresponds to an $M$ resonant multi-soliton (1.20) and $F=\log$ if and only if $\Theta$ satisfies $H\left(\Theta\right)=Ai\left(\Theta\right)=0$ and $\Theta W_{y}\left(\Theta\right)=\Theta W_{x}\left(\Theta\right)$ has a unique value in $\mathcal{W}_{\frac{1}{2}M\left(M-1\right)}$.
@@
The key in the proof of Theorem 1.11 is the property that $\Theta\in\mathcal{W}_{M}$ implies $\Theta W_{y}\left(\Theta\right)\in\mathcal{W}_{\frac{1}{2}M\left(M-1\right)}$ (Lemma 4.7). This property allows us to estimate the size of the cone representing the image of $\mathcal{W}_{M}$ under the nonlinear mapping $\Theta W_{y}\left(\Theta\right)$. Then one has to establish a sort of uniqueness in the representation of $\Theta$, which in the case $M=1$ is easy to obtain (see Theorem 1.8), but in the general case it is not known to hold in generality. Under the additional prescribed data at $x=y=0$, then uniqueness is recovered and Theorem 1.11 establishes the equivalence between resonant multi-solitons and Airy-Heat type phases with finite dimensional Wronskians.
@@
Notice that Theorem 1.13 does not assume that $\Theta=\hbox{Wr}\left[\Theta_{1},\Theta_{2}\right]$ (standard Wronskian of $\Theta_{1}$ and $\Theta_{2}$), with $H\left(\Theta_{i}\right)=Ai\left(\Theta_{i}\right)=0$ and $\Theta_{1},\Theta_{2}\in\mathcal{W}_{2}$. This is the standard and well-known definition of the 2-soliton that assumes the key Wronskian substructure. Here we lift that condition and only ask conditions on $\Theta$ itself. The Wronskian structure is recovered from the proof.
@@
Theorem 1.13 can be recast as follows. It is easy to check that 2-solitons solve $\mathcal{T}\left(\Theta\right)=0$ (Corollary 4.10). However, this equation has plenty of additional, more complicated solutions [22], and a suitable characterization of the 2-soliton subclass is desirable. In that sense, Theorem 1.13 and the $\mathcal{W}_{n}$ structure give a precise equivalence that separates 2-solitons of other more complex KP solutions.
@@ 1.4. Further results
$Ai(\Theta)=W_{1}(\Theta)=W_{x_{j}}^{F}(\Theta)=0,\quad j=2,\ldots,d,$ (1.23)
@@
$W_{1}(\Theta):=(\partial_{x_{1}}^{2}\Theta)^{2}-\partial_{x_{1}}\Theta\partial_{x_{1}}^{4}\Theta,\quad W_{x_{j}}^{F}(\Theta):=\partial_{x_{j}}^{2}\Theta-F^{\prime}(\Theta)(\partial_{x_{j}}\Theta)^{2}.$
@@
$\left(-4u_{t}+\partial_{x_{1}}^{3}u+6u^{2}\partial_{x_{1}}u\right)+\partial_{x_{1}}(\Delta_{c}u)=0,$ (1.24)
@@
Let $u=2\partial_{x_{1}}F(\Theta)$ be a smooth solution of (1.24) with smooth profile $F:\mathbb{R}\to\mathbb{R}$ satisfying $F\left(0\right)=0$ and $F$ strictly increasing in $\mathbb{R}$. Then a nontrivial $\Theta$ is a mKdV soliton (B.11) and $F=2\arctan$, and only if
$Ai(\Theta)=W(\Theta)=\Lambda_{x_{j}}^{F}(\Theta)=0,\quad j=2,\ldots,d,$ (1.25)
@@
$W\left(\Theta\right):=\Theta_{xx}^{2}-\Theta_{x}\Theta_{xx},\quad\Lambda_{x_{j}}^{F}(\Theta):=\Theta_{x_{j}x_{j}}F^{\prime}(\Theta)+\Theta_{x_{j}}^{2}F^{\prime\prime}(\Theta).$
@@ 2.2. Kernel in the solitonic representation
As a consequence of the previous result, if $F=\log$, any phase that can be expressed in the form $\Theta=\exp\left(f\left(t,x,y\right)\right)$ with $f\left(t,x,y\right)$ any smooth linear affine function in the $x$ variable gives a trivial solution. This is the kernel of the operator $\partial^{2}_{x}\log$ which permits to construct the KP multi-soliton solutions, based on this first seed.
@@ 2.3. Quick review of the simplest KP solitons
Now consider the case of resonant solitons (1.15) with $M=3$. As in the previous case, it is possible to determine the dominant exponentials and analyze the structure of the solution in the $xy$-plane. Indeed, the line-soliton at $y$ positive, corresponding to the $\left[1,3\right]$-soliton, is located on the phase transition $x+cy=constant$ with direction parameter $c=k_{1}+k_{3}$. In the same way, the line-soliton located at $y$ negative, corresponding to the $\left[1,2\right]$-soliton and $\left[2,3\right]$-soliton are located over their respective phase transitions with direction parameter $c=k_{1}+k_{2}$ and $c=k_{2}+k_{3}$, respectively. The resonance condition of these three line-solitons is given by
@@ 2. $(ii)$
It can be verified the following isomorphism: $\operatorname{Gr}\left(N,M\right)\cong\operatorname{GL}_{N}\left(\mathbb{R}\right)\setminus M_{N\times M}\left(\mathbb{R}\right)$, where $\operatorname{GL}_{N}\left(\mathbb{R}\right)$ consists of those matrices with dimension $n\times n$, real coefficients and whose determinant is non-zero.
Let $\left\{\Theta_{i}:i=1,\ldots,N\right\}$ be linearly independent solutions of $Ai\left(\Theta\right)=H\left(\Theta\right)=0$. Let $\Theta:=\operatorname{Wr}\left(\Theta_{1},\ldots,\Theta_{N}\right)$ be the Wronskian of the functions $\Theta_{i}$ with respect to the $x$-variable (usually called a $\tau$-function). It is well-known and not difficult to see (see [22]) that $u\left(t,x,y\right)=2\partial^{2}_{x}\ln\left(\Theta\left(t,x,y\right)\right)$ satisfies the KP equation. A particular choice for $\Theta_{i}$ is given by
$\Theta_{i}\left(t,x,y\right)=\sum^{N}_{j=1}a_{ij}\exp\left(\theta_{j}\left(t,x,y\right)\right),\quad\hbox{with}\quad\theta_{j}=k_{j}x+k^{2}_{j}y+k^{3}_{j}t,$

### v2

@@ Abstract.
G.C. was partially funded by Simons Foundation MP-TSM-00002258.
@@ 1.1. Setting of the problem
where $u=u\left(t,x,y\right)\in\mathbb{R}$ is the unknown. The KP equations are canonical integrable models in 2D and were first introduced by Kadomtsev and Petviashvili in 1970 [14] for modeling “long and weakly nonlinear waves” propagating essentially along the $x$ direction, but with a small dependence on the variable $y$. A rigorous derivation of both models from the Boussinesq system was obtained by Lannes and Lannes-Saut [24, 25].
@@
    The third line has a complex structure represented by a Heat type term defined by $H\left(\Theta\right):=\Theta_{y}-\Theta_{xx}$. Notice that $H$ is an operator in the variables $y$ and $x$ only.
@@
Later we will prove that under $H\left(\Theta\right)=0$, one has $W_{x}^{F}\left(\Theta\right)-W_{y}^{F}\left(\Theta\right)=W_{x}\left(\Theta\right)-W_{y}\left(\Theta\right)=0$, namely one can assume that $F=\log$ in (1.11), leading to the natural definitions in (1.6). In that sense, null Wronskians are naturally related to the Heat condition $H\left(\Theta\right)=0$, however, the equivalence will not be as exact as one would prefer.
@@ 1.2. KP solitons
Now we recall the KP 2-soliton. In this case $\Theta=\hbox{Wr}\left(\Theta_{1},\Theta_{2}\right)$, where Wr is the Wronskian of two functions $\Theta_{1}=\exp\left(\theta_{1}\right)+\exp\left(\theta_{2}\right)$ and $\Theta_{2}=\exp\left(\theta_{3}\right)+\exp\left(\theta_{4}\right)$ being 1-soliton phases. Calculating the phase $\Theta$, one obtains the classical formula
@@ 1.3. Main Results
This paper represents a departure from previous approaches, in a sense to be explained below. We adopt the perspective that each KP soliton should adhere to a specific “variational” characterization, manifested in their critical points of a suitable nonlinear functional. Our primary aim is to offer clear and straightforward characterizations of the most distinct KP solitons using simple “trapping” functionals that exhibit rigidity properties. This represents an initial step in the direction previously outlined. We believe that this concept holds promising potential for applications not only in elucidating more complex KP solutions but also in other related dispersive models.
@@
Let $u$ be a smooth solution to KP (1.1) of the form (1.2), with a smooth profile $F$ such that $F\left(1\right)=0$, $F^{\prime}\left(1\right)=1$, $F^{\prime\prime}\left(1\right)=-1$, and $F^{\prime\prime\prime}\left(1\right)=2$. Then $u$ is a KdV soliton and $F=\log$ if and only if $H\left(\Theta\right)=Ai\left(\Theta\right)=W_{y}\left(\Theta\right)=0.$
@@
Assume (1.17). Let $u$ be a nontrivial KP solution of the form (1.2) such that $\Theta>0$ is a solution to $H\left(\Theta\right)=Ai\left(\Theta\right)=W_{y}\left(\Theta\right)=0.$ Then $F=\log$ and $u=Q_{k}$ in (1.14) for some $k\in\mathbb{R}-\left\{0\right\}$.
@@
It is worth to mention that the conditions $H\left(\Theta\right)=Ai\left(\Theta\right)=0$ do not ensure the validity of Theorem 1.8. Indeed, it will be proved that the class of phases satisfying these two conditions is large enough to contain many multi-soliton solutions, such as the $Y$ structure defined below, which is not a line soliton. Consequently, (1.18) is just a necessary condition.
@@
Theorem 1.8 puts in evidence an intriguing new mathematical structure, a natural finite-dimensional cone in the variable $y$.
@@
The space $\mathcal{W}_{n}$ has interesting properties, in particular its behavior under the nonlinear mapping $\Theta W_{y}\left(\Theta\right)$ is key to understand complex multi-soliton structures. First of all, $\Theta\in\mathcal{W}_{M}$ implies that $\Theta W_{y}\left(\Theta\right)\in\mathcal{W}_{\frac{1}{2}M\left(M-1\right)}$ (Lemma 4.7). Second, there is a natural “kernel” given by the function $\exp\left(ky\right)$, $k$ arbitrary: one has $\exp\left(ky\right)W_{y}\left(\exp\left(ky\right)\right)=0$ for any $k=k\left(t,x\right)$. Additionally, under the gauge $\mathcal{W}_{M}\ni\Theta\longmapsto\Theta_{k}:=\exp\left(ky\right)\Theta$, one has $\Theta_{k}W_{y}\left(\Theta_{k}\right)=\exp\left(2ky\right)\Theta W_{y}\left(\Theta\right)\in\mathcal{W}_{\frac{1}{2}M(M-1)}$, revealing that there is no unique nontrivial solution to the set inclusion $\Theta W_{y}\left(\Theta\right)\in\mathcal{W}_{\frac{1}{2}M\left(M-1\right)}$ for $\Theta\in\mathcal{W}_{M}$. One way to repair this gauge freedom is to ask for $\Theta\left(y=0\right)$ and $\Theta_{y}\left(y=0\right)$ uniquely defined, as it is done below. For natural reasons, we will work in a slightly larger class of resonant phases than (1.15), given by
$\Theta\left(t,x,y\right)=\sum^{M}_{i=1}\left(a_{i,1}\exp\left(-k_{i}x+k^{2}_{i}y-k^{3}_{i}t\right)+a_{i,2}\exp\left(k_{i}x+k^{2}_{i}y+k^{3}_{i}t\right)\right),$ (1.20)
with coefficients $k_{1}<k_{2}<\cdots<k_{M}$ and $a_{i,j}\geq 0$.
@@
Assume that the smooth profile $F$ satisfies (1.17). Then $\Theta$ corresponds to an $M$ resonant multi-soliton (1.20) and $F=\log$ if and only if $\Theta$ satisfies $H\left(\Theta\right)=Ai\left(\Theta\right)=0$ and $\Theta W_{y}\left(\Theta\right)=\Theta W_{x}\left(\Theta\right)$ has a unique value in $\mathcal{W}_{\frac{1}{2}M\left(M-1\right)}$.
@@
The key in the proof of Theorem 1.11 is the property that $\Theta\in\mathcal{W}_{M}$ implies $\Theta W_{y}\left(\Theta\right)\in\mathcal{W}_{\frac{1}{2}M\left(M-1\right)}$ (Lemma 4.7). This property allows us to estimate the size of the cone representing the image of $\mathcal{W}_{M}$ under the nonlinear mapping $\Theta W_{y}\left(\Theta\right)$. Then one has to establish a sort of uniqueness in the representation of $\Theta$, which in the case $M=1$ is easy to obtain (see Theorem 1.8), but it is not known if it is maintained in general. Under the additional prescribed data at $x=y=0$, uniqueness is recovered and Theorem 1.11 establishes the equivalence between resonant multi-solitons and Airy-Heat type phases with finite dimensional Wronskians.
@@
Notice that Theorem 1.13 does not assume that $\Theta=\hbox{Wr}\left(\Theta_{1},\Theta_{2}\right)$ (standard Wronskian of $\Theta_{1}$ and $\Theta_{2}$), with $H\left(\Theta_{i}\right)=Ai\left(\Theta_{i}\right)=0$ and $\Theta_{1},\Theta_{2}\in\mathcal{W}_{2}$. This is the standard and well-known definition of the 2-soliton that assumes the key Wronskian substructure. Here we lift that condition and only ask conditions on $\Theta$ itself. The Wronskian structure is recovered from the proof.
@@
Theorem 1.13 can be recast as follows. It is easy to check that 2-solitons (1.16) solve $\mathcal{T}\left(\Theta\right)=0$ when $F=\log$ (Corollary 4.10). However, this equation has plenty of additional, more complicated solutions [22], and a suitable characterization of the 2-soliton subclass is desirable. In that sense, Theorem 1.13 and the $\mathcal{W}_{n}$ structure give a precise equivalence that separates 2-solitons of other more complex KP solutions.
@@ 1.4. Further results
$Ai\left(\Theta\right)=W_{1}\left(\Theta\right)=W_{x_{j}}^{F}\left(\Theta\right)=0,\quad j=2,\ldots,d,$ (1.23)
@@
$W_{1}\left(\Theta\right):=\left(\partial_{x_{1}}^{2}\Theta\right)^{2}-\partial_{x_{1}}\Theta\partial_{x_{1}}^{4}\Theta,\quad W_{x_{j}}^{F}\left(\Theta\right):=\partial_{x_{j}}^{2}\Theta-F^{\prime}\left(\Theta\right)\left(\partial_{x_{j}}\Theta\right)^{2}.$
@@
$\left(-4u_{t}+\partial_{x_{1}}^{3}u+6u^{2}\partial_{x_{1}}u\right)+\partial_{x_{1}}\left(\Delta_{c}u\right)=0,$ (1.24)
@@
Let $u=2\partial_{x_{1}}F\left(\Theta\right)$ be a smooth solution of (1.24) with smooth profile $F:\mathbb{R}\to\mathbb{R}$ satisfying $F\left(0\right)=0$ and $F$ strictly increasing in $\mathbb{R}$. Then a nontrivial $\Theta$ is a mKdV soliton (B.11) and $F=2\arctan$, if and only if
$Ai\left(\Theta\right)=W\left(\Theta\right)=\Lambda_{x_{j}}^{F}\left(\Theta\right)=0,\quad j=2,\ldots,d,$ (1.25)
@@
$W\left(\Theta\right):=\Theta_{xx}^{2}-\Theta_{x}\Theta_{xx},\quad\Lambda_{x_{j}}^{F}\left(\Theta\right):=\Theta_{x_{j}x_{j}}F^{\prime}\left(\Theta\right)+\Theta_{x_{j}}^{2}F^{\prime\prime}\left(\Theta\right).$
@@ 2.2. Kernel in the solitonic representation
As a consequence of the previous result, if $F=\log$, any phase that can be expressed in the form $\Theta=\exp\left(f\left(t,x,y\right)\right)$ with $f\left(t,x,y\right)$ any smooth linear affine function in the variable $x$ gives a trivial solution. This is the kernel of the operator $\partial^{2}_{x}\log$ which permits to construct the KP multi-soliton solutions, based on this first seed.
@@ 2.3. Quick review of the simplest KP solitons
Now consider the case of resonant solitons (1.15) with $M=3$. As in the previous case, it is possible to determine the dominant exponentials and analyze the structure of the solution in the $xy$-plane. Indeed, the line-soliton at positive $y$, corresponding to the $\left[1,3\right]$-soliton, is located on the phase transition $x+cy=constant$ with direction parameter $c=k_{1}+k_{3}$. In the same way, the line-soliton located at $y$ negative, corresponding to the $\left[1,2\right]$-soliton and $\left[2,3\right]$-soliton are located over their respective phase transitions with direction parameter $c=k_{1}+k_{2}$ and $c=k_{2}+k_{3}$, respectively. The resonance condition of these three line-solitons is given by
@@ 2. $(ii)$
It can be verified the following isomorphism: $\operatorname{Gr}\left(N,M\right)\cong\operatorname{GL}_{N}\left(\mathbb{R}\right)\setminus M_{N\times M}\left(\mathbb{R}\right)$, where $\operatorname{GL}_{N}\left(\mathbb{R}\right)$ consists of those matrices with dimension $n\times n$, real coefficients and whose determinant is non-zero, and $M_{N\times M}(\mathbb{R})$ is the set of all $N\times M$ full-rank matrices.
Let $\left\{\Theta_{i}:i=1,\ldots,N\right\}$ be linearly independent solutions of $Ai\left(\Theta\right)=H\left(\Theta\right)=0$. Let $\Theta:=\operatorname{Wr}\left(\Theta_{1},\ldots,\Theta_{N}\right)$ be the Wronskian of the functions $\Theta_{i}$ with respect to the variable $x$ (usually called a $\tau$-function). It is well-known and not difficult to see (see [22]) that $u\left(t,x,y\right)=2\partial^{2}_{x}\log\left(\Theta\left(t,x,y\right)\right)$ satisfies the KP equation. A particular choice for $\Theta_{i}$ is given by
$\Theta_{i}\left(t,x,y\right)=\sum^{M}_{j=1}a_{ij}\exp\left(\theta_{j}\left(t,x,y\right)\right),\quad\hbox{with}\quad\theta_{j}=k_{j}x+k^{2}_{j}y+k^{3}_{j}t,$